# In Praise of Stubbornness: The Case for Cognitive-Dissonance-Aware Knowledge Updates in LLMs

### v1

@@ Appendix A Extended Related work
iCaRL (Rebuffi et al., 2017) Class-incremental Replay Task-Agnostic Fixed Fixed No Rehearsal
EWC (Kirkpatrick et al., 2017b) Task-incremental None Task-Aware Selective Fixed No Regularization
Progressive Nets (Rusu et al., 2016) Task-incremental None Task-Aware Fixed Expanding No New Subnetworks
DEN (Yoon et al., 2017) Task-incremental None Task-Aware Selective Expanding No Selective Expansion
GEM (Lopez-Paz & Ranzato, 2017) Task-incremental Replay Task-Aware Constrained Fixed No Constrained Optimization
ROME (De Cao et al., 2021) Fact-incremental None Fact-Aware Localized Fixed No Rank-One Update
OWM (Zeng et al., 2019) Task-incremental None Task-Aware Orthogonal Fixed No Orthogonal Projection
PackNet (Mallya & Lazebnik, 2018) Task-incremental None Task-Aware Selective Fixed No Weight Masking
HAT (Serra et al., 2018) Task-incremental None Task-Aware Selective Fixed No Attention Masking
MALMEN (Tan et al., 2023) Fact-incremental None Fact-Aware Localized Fixed No Parameter Shift Aggregation
EditAnalysis (Li et al., 2023) Fact-incremental None Fact-Aware Analysis Fixed No Consistency Analysis
D4S (Huang et al., 2024) Fact-incremental O(1) Fact-Aware Regulated Fixed No Layer-Norm Control
Global Prototypes (Bai et al., 2024) Task/Class-incremental None Task-Agnostic Selective Fixed No Global Prototype Alignment
NTE (Benjamin et al., 2024) Task-incremental None Task-Agnostic Selective Fixed No Bayesian Ensemble
UPGD (Elsayed & Mahmood, 2024) Task-incremental None Task-Agnostic Selective Fixed No Utility-Gated Updates
(Hiratani, 2024) Task-incremental None Task-Aware Selective Fixed No Fisher Information
CLAP (Jha et al., 2024) Class-incremental None Task-Aware Selective Fixed No Probabilistic Adaptation
VQ-Prompt (Jiao et al., 2024) Class-incremental None Task-Agnostic Fixed Fixed No Discrete Prompt Selection
IsCiL (Lee et al., 2024) Task-incremental None Task-Aware Selective Fixed No Skill-based Adaptation
BGS (Lee et al., ) Task/Domain/Class-incremental Replay Task-Aware Selective Fixed Yes Bias-Aware Update
SLM (Peng et al., 2024) Task-incremental None Auto-detected Selective Fixed No Vector Space Retrieval
Train-Attention (Seo et al., 2024) Knowledge-incremental None Task-Agnostic Selective Fixed No Token-Weighted Update
Refresh Learning (Wang et al., 2024) Task/Class-incremental Optional Task-Aware Selective Fixed No Unlearn-Relearn
RAIL (Xu et al., 2024) Cross-domain-incremental None Task-Agnostic Selective Fixed No Regression-based Update
SAFE (Zhao et al., 2024) Class-incremental None Task-Agnostic Selective Fixed No Dual Parameter-Efficient Tuning
This paper Fact-incremental None Conflict-Aware Selective Fixed Yes Neuron-Specific Update
@@ A.1 Continual learning
Continual Learning (CL) methods enable models to learn new tasks without catastrophically forgetting previously mastered ones (Kirkpatrick et al., 2017b). These approaches fall into three main families: memory-based methods using exemplar buffers (Rebuffi et al., 2017), knowledge distillation techniques that transfer information across model versions (Lopez-Paz & Ranzato, 2017), and regularization-based methods that constrain weight updates (Kirkpatrick et al., 2017b). To ease the understanding of this landscape, we build a taxonomy that characterizes approaches by their incremental type (task, class, or fact-based), memory requirements, update mechanisms, and architectural constraints (Tab. 1). This taxonomy reveals how our work is different from existing continual learning attempts: while existing methods focus on preserving knowledge across distinct tasks, none explicitly address the detection and handling of conflicting information - a key capability in human cognition that our work empirically investigates.
One of the closest old approaches is deep mind’s EWC (Kirkpatrick et al., 2017b), a method designed to mitigate catastrophic forgetting in neural networks trained sequentially on distinct tasks. The core idea is to protect the most important weights (or neurons) for previously learned tasks during the training of new tasks. EWC identifies these important weights by calculating the Fisher Information Matrix during or after the training of a task, which estimates how sensitive each weight is to the task’s performance. Weights that significantly impact the output for a given task are marked as important. A quadratic penalty is then applied during future learning, constraining these weights to remain close to their values from the previous task. This ensures that knowledge from earlier tasks is preserved while still allowing the model to adapt to new tasks. However, EWC is less suitable for LLMs, which do not have clearly defined tasks when it comes to knowledge ingestion (probably different for other types of skills). EWC’s effectiveness relies on distinct task boundaries and the ability to compute task-specific importance for weights, which is feasible in scenarios with well-defined tasks, such as classification or reinforcement learning. In LLMs, where learning spans a wide range of topics and linguistic structures without clear task delineation, it’s challenging to apply EWC’s task-based strategy. The model would struggle to assign specific neurons or weights to individual tasks or concepts, making it difficult to protect task-specific knowledge without hindering the model’s overall generalization ability across a diverse dataset.
@@
Bai et al. (2024) introduce a novel approach to continual learning that leverages global prototypes to mitigate catastrophic forgetting in neural networks. Their key insight is that maintaining stable connections between task-specific representations and pre-learned, general-purpose token embeddings (which serve as global prototypes) can significantly reduce forgetting without requiring explicit replay mechanisms. Through empirical validation on both task-incremental and class-incremental NLP scenarios, they demonstrate that models preserving strong connections to these global prototypes exhibit enhanced stability. While their work shares our goal of preserving knowledge during updates, it differs fundamentally in its approach and granularity: where they focus on task-level knowledge preservation through architectural mechanisms, our work addresses the more specific challenge of managing contradictory factual updates through cognitive-inspired conflict detection. Their finding that stable reference points aid knowledge retention is conceptually relevant to our work, though our results suggest that such architectural approaches alone may be insufficient when handling explicitly contradictory information, where more sophisticated cognitive mechanisms become necessary.
@@
Benjamin et al. (2024) proposed an elegant theoretical framework that interprets neural networks as Bayesian ensembles of classifiers. Their key insight is that a neural network with N parameters can be viewed as a weighted ensemble of N classifiers in the lazy regime, where the classifiers remain fixed throughout learning. This interpretation reveals that a properly designed posterior update rule, resembling SGD without momentum, can enable continual learning without forgetting - notably, they prove that momentum actually exacerbates forgetting. While their work focuses on preserving all knowledge in task-incremental learning, our paper specifically examines cases where knowledge needs to be deliberately updated or overridden. Their key contribution is showing that catastrophic forgetting is linked to the transition from lazy to rich regimes in neural networks, providing both a theoretical explanation for why larger models are more robust to forgetting and a biologically-inspired mechanism for knowledge preservation that perhaps complements our cognitive-based approach.
@@
Elsayed & Mahmood (2024) propose UPGD (Utility-based Perturbed Gradient Descent), a novel approach targeting both catastrophic forgetting and loss of plasticity in streaming learning scenarios. Their method protects useful network units while maintaining plasticity in less-used ones through utility-gated gradient updates and perturbations. Unlike previous approaches requiring task boundaries or memory buffers, UPGD operates in a challenging streaming setting with continuous non-stationarity. Using their newly introduced direct plasticity metric, they demonstrate UPGD’s ability to maintain performance levels that surpass or match existing methods. This work complements our investigation by providing evidence that selective neuronal updates based on utility metrics can effectively balance stability and plasticity, though in a task-learning rather than knowledge-updating context.
Hiratani (2024) analyze how task similarity affects continual learning through a novel theoretical framework combining teacher-student models with latent structure. Their key insight is that high input feature similarity coupled with low readout similarity leads to catastrophic outcomes in both knowledge transfer and retention, even when tasks are positively correlated. They demonstrate that weight regularization in the Fisher information metric robustly helps retention regardless of task similarity, while common approaches like activity gating improve retention at the cost of transfer performance. Their theoretical predictions are validated on permuted MNIST tasks with latent variables.
@@
Jha et al. (2024) propose a probabilistic approach to continual learning for vision-language models, specifically focusing on CLIP adaptation. Their method, CLAP, introduces visual-guided attention and task-specific probabilistic adapters to model the distribution of text features, while leveraging CLIP’s pre-trained knowledge for initialization and regularization. This work demonstrates that probabilistic modeling can significantly reduce catastrophic forgetting in class-incremental learning scenarios, achieving state-of-the-art performance across multiple benchmarks.
@@
Jiao et al. (2024) propose VQ-Prompt, a novel prompt-based continual learning framework that addresses class-incremental learning with pretrained vision transformers. Their key innovation is incorporating vector quantization into prompt selection, enabling end-to-end optimization of discrete prompts with task loss while maintaining effective knowledge abstraction. This contrasts with our cognitive-dissonance aware approach, as they focus on task adaptation through prompt engineering rather than explicit conflict detection. Their empirical results on ImageNet-R and CIFAR-100 demonstrate superior performance compared to existing prompt-based methods, suggesting the effectiveness of discrete knowledge representation in continual learning.
@@
(Lee et al., 2024) propose IsCiL, a framework for continual imitation learning that uses retrievable skills and adapter-based architecture to enable efficient knowledge sharing across tasks. Unlike traditional approaches that isolate task-specific parameters, IsCiL introduces a prototype-based skill retrieval mechanism that allows selective reuse of previously learned skills for new tasks. While focused primarily on motor skills rather than resolving knowledge contradictions, their empirical results show that this selective adaptation approach significantly improves sample efficiency and reduces catastrophic forgetting compared to other adapter-based methods, particularly in scenarios with incomplete demonstrations.
@@
Peng et al. (2024) proposed a continual learning approach that automates task selection through vector space retrieval, eliminating the need for explicit task IDs, experience replay, or optimization constraints. Their method, Scalable Language Model (SLM), combines Joint Adaptive Re-parameterization with dynamic knowledge retrieval to automatically identify relevant parameters for each input, enabling task-agnostic updates. While achieving state-of-the-art results across diverse tasks and model scales (BERT, T5, LLaMA-2), their key contribution is demonstrating that automatic task identification and parameter selection can enable continual learning without requiring explicit task boundaries or memory buffers.
@@
Seo et al. (2024) presented Train-Attention, an interesting meta-learning approach for continual knowledge learning (CKL) in LLMs that predicts and applies weights to tokens based on their usefulness for future tasks. Unlike previous approaches that uniformly update all parameters, their method enables targeted knowledge updates by learning which tokens are most important to focus on. Through experiments on LAMA-CKL and TemporalWiki benchmarks, they show that selective token-weighted learning significantly reduces catastrophic forgetting while improving learning speed. The work somewhat complements our cognitive-inspired approach, and demonstrates the benefits of selective attention, but it does not explicitly address the handling of contradictory information.
@@
Wang et al. (2024) proposed a unified framework for continual learning that reveals common mathematical structures across seemingly distinct approaches (regularization-based, Bayesian-based, and memory-replay). Building on this unification, they introduce “refresh learning” - a plug-in mechanism that first unlearns current data before relearning it, inspired by the beneficial role of forgetting in human cognition. Their work primarily focuses on task-incremental and class-incremental scenarios, demonstrating improved accuracy across CIFAR and Tiny-ImageNet benchmarks. While their approach differs from our fact-level knowledge updates in LLMs, their findings about selective forgetting complement our observations about cognitive-inspired update mechanisms. Their theoretical analysis showing that refresh learning improves the flatness of the loss landscape offers an interesting perspective on how controlled forgetting might benefit knowledge retention in neural networks.
@@
Xu et al. (2024) propose a cross-domain task-agnostic incremental learning framework (X-TAIL) for vision-language models, focusing on the challenge of preserving both incrementally learned knowledge and zero-shot abilities. Their approach, RAIL, uses recursive ridge regression with non-linear projections to adapt to new domains without catastrophic forgetting. Unlike previous work requiring domain identity hints or reference datasets, RAIL can classify images across both seen and unseen domains without domain hints, demonstrating superior performance in both discriminative ability and knowledge preservation. While their work advances the technical aspects of continual learning, it differs from our cognitive-inspired investigation as it doesn’t address the fundamental challenge of detecting and resolving conflicting knowledge, instead focusing on domain adaptation without explicit conflict awareness.
@@
Zhao et al. (2024) propose a class-incremental learning framework for pre-trained vision models that balances stability and plasticity through two complementary parameter-efficient tuning mechanisms. Their SAFE approach first inherits generalizability from pre-trained models via a “slow learner” that captures transferable knowledge in the first session, then maintains plasticity through a “fast learner” that continuously adapts to new classes while resisting catastrophic forgetting. While focused on vision tasks rather than language models, their dual-speed learning strategy presents interesting parallels to our cognitive-inspired approach – particularly in how both works identify the importance of selective plasticity and the distinction between stable (“stubborn”) and adaptable (“plastic”) parameters. However, SAFE doesn’t address the fundamental challenge of detecting and handling contradictory information that we identify as crucial for true cognitive-inspired learning.

### v2

@@ Appendix A Extended Related work
iCaRL [Rebuffi et al., 2017] Class-incremental Replay Task-Agnostic Fixed Fixed No Rehearsal
EWC [Kirkpatrick et al., 2017] Task-incremental None Task-Aware Selective Fixed No Regularization
Progressive Nets [Rusu et al., 2016] Task-incremental None Task-Aware Fixed Expanding No New Subnetworks
DEN [Yoon et al., 2017] Task-incremental None Task-Aware Selective Expanding No Selective Expansion
GEM [Lopez-Paz and Ranzato, 2017] Task-incremental Replay Task-Aware Constrained Fixed No Constrained Optimization
ROME [De Cao et al., 2021] Fact-incremental None Fact-Aware Localized Fixed NA Rank-One Update
OWM [Zeng et al., 2019] Task-incremental None Task-Aware Orthogonal Fixed No Orthogonal Projection
PackNet [Mallya and Lazebnik, 2018] Task-incremental None Task-Aware Selective Fixed No Weight Masking
HAT [Serra et al., 2018] Task-incremental None Task-Aware Selective Fixed No Attention Masking
MALMEN [Tan et al., 2023] Fact-incremental None Fact-Aware Localized Fixed NA Parameter Shift Aggregation
EditAnalysis [Li et al., 2023] Fact-incremental None Fact-Aware Analysis Fixed NA Consistency Analysis
D4S [Huang et al., 2024] Fact-incremental O(1) Fact-Aware Regulated Fixed NA Layer-Norm Control
Global Prototypes [Bai et al., 2024] Task/Class-incremental None Task-Agnostic Selective Fixed No Global Prototype Alignment
NTE [Benjamin et al., 2024] Task-incremental None Task-Agnostic Selective Fixed No Bayesian Ensemble
UPGD [Elsayed and Mahmood, 2024] Task-incremental None Task-Agnostic Selective Fixed No Utility-Gated Updates
CLAP [Jha et al., 2024] Class-incremental None Task-Aware Selective Fixed No Probabilistic Adaptation
VQ-Prompt [Jiao et al., 2024] Class-incremental None Task-Agnostic Fixed Fixed No Discrete Prompt Selection
IsCiL [Lee et al., 2024] Task-incremental None Task-Aware Selective Fixed No Skill-based Adaptation
BGS [Lee et al., ] Task/Domain/Class-incremental Replay Task-Aware Selective Fixed Yes Bias-Aware Update
SLM [Peng et al., 2024] Task-incremental None Auto-detected Selective Fixed No Vector Space Retrieval
Train-Attention [Seo et al., 2024] Knowledge-incremental None Task-Agnostic Selective Fixed No Token-Weighted Update
Refresh Learning [Wang et al., 2024] Task/Class-incremental Optional Task-Aware Selective Fixed No Unlearn-Relearn
RAIL [Xu et al., 2024] Cross-domain-incremental None Task-Agnostic Selective Fixed No Regression-based Update
SAFE [Zhao et al., 2024] Class-incremental None Task-Agnostic Selective Fixed No Dual Parameter-Efficient Tuning
@@ A.1 Continual learning
Continual Learning (CL) methods enable models to learn new tasks without catastrophically forgetting previously mastered ones [Kirkpatrick et al., 2017]. These approaches fall into three main families: memory-based methods using exemplar buffers [Rebuffi et al., 2017], knowledge distillation techniques that transfer information across model versions [Lopez-Paz and Ranzato, 2017], and regularization-based methods that constrain weight updates [Kirkpatrick et al., 2017]. To ease the understanding of this landscape, we build a taxonomy that characterizes approaches by their incremental type (task, class, or fact-based), memory requirements, update mechanisms, and architectural constraints (Tab. 1). This taxonomy reveals how our work is different from existing continual learning attempts: while existing methods focus on preserving knowledge across distinct tasks, none explicitly address the detection and handling of conflicting information - a key capability in human cognition that our work empirically investigates.
One of the closest old approaches is deep mind’s EWC [Kirkpatrick et al., 2017], a method designed to mitigate catastrophic forgetting in neural networks trained sequentially on distinct tasks. The core idea is to protect the most important weights (or neurons) for previously learned tasks during the training of new tasks. EWC identifies these important weights by calculating the Fisher Information Matrix during or after the training of a task, which estimates how sensitive each weight is to the task’s performance. Weights that significantly impact the output for a given task are marked as important. A quadratic penalty is then applied during future learning, constraining these weights to remain close to their values from the previous task. This ensures that knowledge from earlier tasks is preserved while still allowing the model to adapt to new tasks. However, EWC is less suitable for LLMs, which do not have clearly defined tasks when it comes to knowledge ingestion (probably different for other types of skills). EWC’s effectiveness relies on distinct task boundaries and the ability to compute task-specific importance for weights, which is feasible in scenarios with well-defined tasks, such as classification or reinforcement learning. In LLMs, where learning spans a wide range of topics and linguistic structures without clear task delineation, it’s challenging to apply EWC’s task-based strategy. The model would struggle to assign specific neurons or weights to individual tasks or concepts, making it difficult to protect task-specific knowledge without hindering the model’s overall generalization ability across a diverse dataset.
@@
Bai et al. [2024] introduce a novel approach to continual learning that leverages global prototypes to mitigate catastrophic forgetting in neural networks. Their key insight is that maintaining stable connections between task-specific representations and pre-learned, general-purpose token embeddings (which serve as global prototypes) can significantly reduce forgetting without requiring explicit replay mechanisms. Through empirical validation on both task-incremental and class-incremental NLP scenarios, they demonstrate that models preserving strong connections to these global prototypes exhibit enhanced stability. While their work shares our goal of preserving knowledge during updates, it differs fundamentally in its approach and granularity: where they focus on task-level knowledge preservation through architectural mechanisms, our work addresses the more specific challenge of managing contradictory factual updates through cognitive-inspired conflict detection. Their finding that stable reference points aid knowledge retention is conceptually relevant to our work, though our results suggest that such architectural approaches alone may be insufficient when handling explicitly contradictory information, where more sophisticated cognitive mechanisms become necessary.
@@
Benjamin et al. [2024] proposed an elegant theoretical framework that interprets neural networks as Bayesian ensembles of classifiers. Their key insight is that a neural network with N parameters can be viewed as a weighted ensemble of N classifiers in the lazy regime, where the classifiers remain fixed throughout learning. This interpretation reveals that a properly designed posterior update rule, resembling SGD without momentum, can enable continual learning without forgetting - notably, they prove that momentum actually exacerbates forgetting. While their work focuses on preserving all knowledge in task-incremental learning, our paper specifically examines cases where knowledge needs to be deliberately updated or overridden. Their key contribution is showing that catastrophic forgetting is linked to the transition from lazy to rich regimes in neural networks, providing both a theoretical explanation for why larger models are more robust to forgetting and a biologically-inspired mechanism for knowledge preservation that perhaps complements our cognitive-based approach.
@@
Elsayed and Mahmood [2024] propose UPGD (Utility-based Perturbed Gradient Descent), a novel approach targeting both catastrophic forgetting and loss of plasticity in streaming learning scenarios. Their method protects useful network units while maintaining plasticity in less-used ones through utility-gated gradient updates and perturbations. Unlike previous approaches requiring task boundaries or memory buffers, UPGD operates in a challenging streaming setting with continuous non-stationarity. Using their newly introduced direct plasticity metric, they demonstrate UPGD’s ability to maintain performance levels that surpass or match existing methods. This work complements our investigation by providing evidence that selective neuronal updates based on utility metrics can effectively balance stability and plasticity, though in a task-learning rather than knowledge-updating context.
@@
Jha et al. [2024] propose a probabilistic approach to continual learning for vision-language models, specifically focusing on CLIP adaptation. Their method, CLAP, introduces visual-guided attention and task-specific probabilistic adapters to model the distribution of text features, while leveraging CLIP’s pre-trained knowledge for initialization and regularization. This work demonstrates that probabilistic modeling can significantly reduce catastrophic forgetting in class-incremental learning scenarios, achieving state-of-the-art performance across multiple benchmarks.
@@
Jiao et al. [2024] propose VQ-Prompt, a novel prompt-based continual learning framework that addresses class-incremental learning with pretrained vision transformers. Their key innovation is incorporating vector quantization into prompt selection, enabling end-to-end optimization of discrete prompts with task loss while maintaining effective knowledge abstraction. This contrasts with our cognitive-dissonance aware approach, as they focus on task adaptation through prompt engineering rather than explicit conflict detection. Their empirical results on ImageNet-R and CIFAR-100 demonstrate superior performance compared to existing prompt-based methods, suggesting the effectiveness of discrete knowledge representation in continual learning.
@@
Lee et al. [2024] propose IsCiL, a framework for continual imitation learning that uses retrievable skills and adapter-based architecture to enable efficient knowledge sharing across tasks. Unlike traditional approaches that isolate task-specific parameters, IsCiL introduces a prototype-based skill retrieval mechanism that allows selective reuse of previously learned skills for new tasks. While focused primarily on motor skills rather than resolving knowledge contradictions, their empirical results show that this selective adaptation approach significantly improves sample efficiency and reduces catastrophic forgetting compared to other adapter-based methods, particularly in scenarios with incomplete demonstrations.
@@
Peng et al. [2024] proposed a continual learning approach that automates task selection through vector space retrieval, eliminating the need for explicit task IDs, experience replay, or optimization constraints. Their method, Scalable Language Model (SLM), combines Joint Adaptive Re-parameterization with dynamic knowledge retrieval to automatically identify relevant parameters for each input, enabling task-agnostic updates. While achieving state-of-the-art results across diverse tasks and model scales (BERT, T5, LLaMA-2), their key contribution is demonstrating that automatic task identification and parameter selection can enable continual learning without requiring explicit task boundaries or memory buffers.
@@
Seo et al. [2024] presented Train-Attention, an interesting meta-learning approach for continual knowledge learning (CKL) in LLMs that predicts and applies weights to tokens based on their usefulness for future tasks. Unlike previous approaches that uniformly update all parameters, their method enables targeted knowledge updates by learning which tokens are most important to focus on. Through experiments on LAMA-CKL and TemporalWiki benchmarks, they show that selective token-weighted learning significantly reduces catastrophic forgetting while improving learning speed. The work somewhat complements our cognitive-inspired approach, and demonstrates the benefits of selective attention, but it does not explicitly address the handling of contradictory information.
@@
Wang et al. [2024] proposed a unified framework for continual learning that reveals common mathematical structures across seemingly distinct approaches (regularization-based, Bayesian-based, and memory-replay). Building on this unification, they introduce “refresh learning" - a plug-in mechanism that first unlearns current data before relearning it, inspired by the beneficial role of forgetting in human cognition. Their work primarily focuses on task-incremental and class-incremental scenarios, demonstrating improved accuracy across CIFAR and Tiny-ImageNet benchmarks. While their approach differs from our fact-level knowledge updates in LLMs, their findings about selective forgetting complement our observations about cognitive-inspired update mechanisms. Their theoretical analysis showing that refresh learning improves the flatness of the loss landscape offers an interesting perspective on how controlled forgetting might benefit knowledge retention in neural networks.
@@
Xu et al. [2024] propose a cross-domain task-agnostic incremental learning framework (X-TAIL) for vision-language models, focusing on the challenge of preserving both incrementally learned knowledge and zero-shot abilities. Their approach, RAIL, uses recursive ridge regression with non-linear projections to adapt to new domains without catastrophic forgetting. Unlike previous work requiring domain identity hints or reference datasets, RAIL can classify images across both seen and unseen domains without domain hints, demonstrating superior performance in both discriminative ability and knowledge preservation. While their work advances the technical aspects of continual learning, it differs from our cognitive-inspired investigation as it doesn’t address the fundamental challenge of detecting and resolving conflicting knowledge, instead focusing on domain adaptation without explicit conflict awareness.
@@
Zhao et al. [2024] propose a class-incremental learning framework for pre-trained vision models that balances stability and plasticity through two complementary parameter-efficient tuning mechanisms. Their SAFE approach first inherits generalizability from pre-trained models via a “slow learner" that captures transferable knowledge in the first session, then maintains plasticity through a “fast learner" that continuously adapts to new classes while resisting catastrophic forgetting. While focused on vision tasks rather than language models, their dual-speed learning strategy presents interesting parallels to our cognitive-inspired approach – particularly in how both works identify the importance of selective plasticity and the distinction between stable (“stubborn") and adaptable (“plastic") parameters. However, SAFE doesn’t address the fundamental challenge of detecting and handling contradictory information that we identify as crucial for true cognitive-inspired learning.